OpenStudy (anonymous):

Prove that (sinx+cosx)(tan^2x+1/tanx)=1/cosx+1/sinx

OpenStudy (anonymous):

my teacher doesnt want us using tan identities so we have to switch all the tans to sinx/cosx first

OpenStudy (anonymous):

so far RS=sinx+cosx/cosxsinx

OpenStudy (cwrw238):

sorry i misinterpreted the question right so tan^2 x = sin^2 x/ cos^2 x and 1/tanx = cos x / sin x

OpenStudy (anonymous):

LS=(sinx+cosx)((sin^2x/cos^2x)+1/(sinx/cosx))

OpenStudy (cwrw238):

that's horrendously complicated are you sure you have copied the question correctly?

OpenStudy (cwrw238):

in your second bracket could it be (tan^x + 1) / tanx )?

OpenStudy (anonymous):

no its tan^2+1/tanx

OpenStudy (anonymous):

sorry tan^2x+1

OpenStudy (anonymous):

i double checked what i typed in, what you see above is the question i see on my paper

OpenStudy (cwrw238):

if that the case then I see a way to prove it LS = ( sinx + cosx) (sec^x / tanx) = (sinx + cosx ( cos x / sin x * 1/cos^2x) = (sin x + cos x) * 1 / sin x cos x = RS

OpenStudy (anonymous):

how did you get (sec^2x/tanx?)

OpenStudy (cwrw238):

tan^2 x + 1 = sec^2x and its divided by tan x also sec^2 x = 1 / cos^2x so sec^2 x / tan x = (1/ cos^2x) * cos x / sinx = 1 /sinx cosx

OpenStudy (cwrw238):